# Energy relaxation in a vacuum capacitor-resistor circuit: measurement of sequential decays with divergent time constants

### 23 Jan 2024 v4

@@ Abstract
The decay of the electrical energy in a resistor-vacuum capacitor circuit is shown to involve multiple sequential relaxation processes, with increasing time constants, separated by transition regions that involve a change in current direction. This is measured using a vacuum capacitor to eliminate the effect of a dielectric between the plates (although ceramic and polypropylene capacitors exhibit similar behavior). A simple phenomenological model accounts for this behavior in spite of the difficulty in applying Maxwell’s equations to such a circuit. These results will lead to a revision of our understanding of the physics of circuits, having particular impact on applications that use capacitors as sensors in collecting precision data (such as found in quantum measurements and dielectric spectroscopy).
@@ I Introduction
To accurately model decay in a vacuum capacitor one must turn to Maxwell’s equations. However, only a handful of solutions that involve steady state behavior have been addressed analytically. Muller (2012) These require a surface charge on the circuit wires to maintain the current as described theoretically Sommerfeld (1952); Heald (1984); Chabay and Sherwood (2019) and utilized experimentally (where the surface charge functions as a Hall probe). Schade et al. (2019) Applying Maxwell’s equations numerically to the relaxation of an RC circuit yields an initial non-exponential decay due to transit time effects. Preyer (2002) However, this calculation does not predict the later sequential decays that are documented below.
@@
There are few first principles analytical models of transient behavior in circuits, even obtaining numerical solutions is challenging. Klee (2020); Muller (2012); Heald (1984) The objective here is to provide a phenomenological model for the observed novel circuit behavior and relate that to a microscopic description of this energy relaxation process. The data are obtained with both different digital voltmeters and capacitors over time constants that vary by five orders of magnitude.
@@ II Results
Figure 1: Current through $R_{1}$ (solid) and $R_{2}$ (dashed) from the circuit in the inset of (a) using a $505$ pF vacuum capacitor with $V=9.4$ is shown in frames (a), (b), and (c). The solid circles in (b) and (d) represent the difference between the currents into and out of the capacitor. The solid circles in (c) represent data for replacing the vacuum capacitor with a $509$ pF ceramic capacitor. All traces are raw data for a single decay. Dips that do not reach the noise baseline (around $\log_{e}|I_{R}|\approx-15$) are data acquisition artifacts.
Figs. 1 (a), (b), and (c) illustrate the decay current $I_{R_{1}}$ through $R_{1}=3930\Omega$ (solid line) and $I_{R_{2}}$ through $R_{2}=3920\Omega$ (dashed line), for the circuit shown in the inset. The vacuum capacitor has $C=505$ pF while $V=9.4$. $M_{1}$ and $M_{2}$ represent the meters that measure the voltages and by Ohm’s law these currents. A mechanical switch disconnects the battery from the RC circuit (with an open switch capacitance of less than $0.1$ pF).
Fig. 1(b) shows an enlarged segment of fig. 1(a). The solid circles represent $I_{R_{1}}+I_{R_{2}}$, indicating that the current entering the capacitor is greater than that leaving it. Fig. 1 (c) illustrates these data on a semi-log plot of the absolute value of the current. The interpretation of such graphs is best demonstrated by comparing fig. 1 (b) with (c). The dips in the semi-log plots correspond to the zero currents in fig. 1(a). After going negative (reversing direction) the current then slowly approaches zero, indicated by a downward slope in the semi-log plot. The straight dot-two-dash line in fig. 1(b) shows the results for the exponential decay with a time constant of $(R_{1}+R_{2})C$. This line is offset along the ordinate for clarity.
In addition to illustrating the vacuum capacitor decay data, the solid circles in frame (c) show the decay data from a $509$ pF ceramic capacitor for the same values of $R_{1}$ and $R_{2}$. These currents from the ceramic capacitor through each resistor are easily separated by the context in which they appear in the figure. This decay behavior appears to be a universal characteristics of capacitors.
@@
The unequal currents into and out of a capacitor result in a net charge on the capacitor. In analogy with the net charge on a hairpin wire this charge plays a role in the decay behavior.
Figure 2: Decay data for the circuit in fig. 1 with $R_{2}=0$, $C=505$ pF, $V=9.4$, and increasing values of $R_{1}$ from top to bottom ($R_{1}=861,~{}7850,~{}27000,~{}63000,$ and $91600~{}\Omega$) plotted vs time constant units in frames (a) and in (b) without the $91600~{}\Omega$ data. The data for $R_{1}=63000~{}\Omega$ in frame (a) is represented by a dashed line. Model results for the $R_{1}=7850~{}\Omega$ decay are represented by solid squares. Frame (c) shows the inverse of the model’s second decay constant (the second time constant) in seconds as a function of $R_{1}$. Frame (d) presents the value of $I_{20}$ as determined from the sum of exponentials model of the data in frames (a) and (b) while the solid line represents the function $0.13/R_{1}$.
Fig. 2 presents decay data for the circuit of fig. 1(a) with the same vacuum capacitor but with $R_{2}=0$ while $R_{1}=861,~{}7850,~{}27000,~{}63000,$ and $91600~{}\Omega$. These are shown with the smallest to largest resistor values from top to bottom at $t/R_{1}C=0$ in fig. 2(a). The same order is shown in fig. 2(b) without the $R_{1}=91600~{}\Omega$ data. Each trace is an average over four decays. A moving average over $20$ data points ($2000$ data points per trace) in frame (b) has been applied to the tails of the data to mitigate overlap of the traces.
Prominent characteristics of the fig. 2 data are: (I) The first relaxation time always matches $1/R_{1}C$ within error while the second, $\beta$ (shown in fig. 2(c)), and third, $\gamma=0.0032\pm.002+(3\pm 4)\times 10^{-8}R_{1}s^{-1}$, are essentially independent of $R_{1}$. (II) Two current reversals occur for each trace (the second dip for the upper trace is not shown since it occurs at $t/\textrm{R}_{1}\textrm{C}=930$). (III) A fit of the time at which the first current reversal occurs yields $T_{1}=1.7\times 10^{-9}R_{1}+2.6\times 10^{-14}R_{1}^{2}$, while that of $T_{2}=0.00040\pm 2\times 10^{-5}s$ is independent of $R_{1}$.
@@ III Phenomenological Model
A model for the $R_{1}=7850~{}\Omega$ data, shown as solid squares in fig. 2 (a) and (b), is given by $I=I_{10}\exp[-\alpha t]-I_{20}\exp[-\beta t]+I_{30}\exp[-\gamma t]]$. This is referred to as the sum of exponentials model. For all of the fits presented here $\alpha=1/\textrm{RC}$, where the values of R and C are determined from measurements of the individual circuit components. Errors in these measurements do not lead to noticeable differences in the model fits shown in the figures.
The parameters of this model for the $R_{1}=7850~{}\Omega$ data are $\alpha=2.27\times 10^{5},\beta=1.2\times 10^{4},\gamma=330$ and $I_{10}=10^{-3},I_{20}=2.5\times 10^{-5},I_{30}=2.8\times 10^{-7}$. This model applied to all the data in fig. 2(a) and (b) results in similar fits to that of the $R_{1}=7850~{}\Omega$ data. The two variable parameters in these fits are $I_{20}$ and $\beta$, a summary of which for the data in fig. 2 is found in (c) and (d).
To measure the decay as the charge distributes over an isolated conductor the single capacitor in fig. 2 is replaced with two in series, as shown in the inset of fig 3 (a) with switches that operate synchronously (within a millisecond while the time constant is $60$ milliseconds). When the switches are closed the isolated right plates of polypropylene capacitors $C_{1}=2.012\mu F$ and $C_{2}=2.014\mu F$ have charge and no charge, respectively. When the switches open the right plates then have a initial net charge which is distributed between them during the decay. Without a switch across $C_{2}$ the net charge on the right plates of $C_{1}$ and $C_{2}$ is initially zero (they are grounded before data collection commences). The response of the circuit for these two cases is shown in figs. 3 (b) and (c). The latter demonstrates enhancement in the non-exponential decay due to this net charge on the right plates of $C_{1}$ and $C_{2}$.
@@
Figure 3: Frames (a), (c), and (d) show meter $M_{2}$ voltage data for the circuit in the inset of frame (a). Voltage data for meter $M_{1}$ is shown in frame (b). The insets in frames (b) and (c) show $C_{dyn}$ for the respective models of the $M_{2}$ data (in microfarads for (b)) while (d) shows the response for different initial voltages. The capacitors are each $2.01\mu F$. For frame (a) through (c) $R=59.5K\Omega$ with $V=9.8$. For frame (d) $R=578K\Omega$ with $V=20,0.74$, and $0.088$ with a microvolt noise baseline and error bars displayed only for a few of the data points.
@@ IV Dynamic Capacitance Model
[…]
@@ V Surface Charge Model
Another modeling approach is to separate the amount of charge on a capacitor plate into two parts: one associated with the energy stored in between the capacitor plates, $Q^{in}$, whose electric field is confined in between the plates, and $Q^{out}$, the charge on the circuit wire and capacitor plate that facilitates the flow of current in the circuit with its electric field not confined. The superposition principle yields a response that is a sum of responses for each type of charge in this separate charge model. Let the $V_{10}\exp[-\alpha t]$ term in the above sum of exponentials model be associated with $Q^{in}$. Let $Q^{out}_{0}$ be the charge required to direct the current from the battery into $R$ before the switch is opened. This is expressed as $Q^{out}_{0}=V^{out}_{0}\tau_{out}/R$, where $\tau_{out}$ is a constant and $V^{out}_{0}$ is the voltage at the plate generated by $Q^{out}_{0}$. For smaller $R$ larger $Q^{out}_{0}$ is required to direct this larger current away from the capacitor plate and into the resistor (similar to the effect that charge on the hairpin wire has on the current flowing through the wire).
@@
The geometry of a circuit increases the difficulty in calculating $Q^{out}$ from Maxwell’s equations. Muller (2012) Yet the data illustrate similar but simple (a sum of exponentials) behavior for dissimilar capacitors. An example is the small ceramic and much larger vacuum capacitor both with the same capacitance but with dissimilar areas generated by the circuit wires and areas of the capacitor plates. In addition, currents due to $Q^{out}$ are measured even after $Q^{in}$ has exponentially decayed over a hundred $\mathrm{RC}$ time constants. During such times $Q^{in}$ is so small that it cannot effect the currents due to $Q^{out}$ that are observed. Another decoupling of $Q^{out}$ occurs between it and the third term in the sum of exponentials model for even longer times, as shown for the $R_{1}=861\Omega$ and $R_{1}=7850\Omega$ decays in fig. 2. This indicates that the third decay is a mode of relaxation different from that of the second decay.
@@ VI Discussion
Constructing a numerical model that predicts only a single exponential decay for an RC circuit from Maxwell’s equations is nontrivial. Preyer (2002) Yet, the data indicating sequential decays with divergent time constants is supported by the simple sum of exponentials model.
@@ VII Conclusions
The decay of the electrical energy in a resistor-vacuum capacitor circuit involves multiple sequential relaxation processes, with increasing time constants, separated by transition regions that involve a change in current direction. This novel behavior differs fundamentally from the single exponential decay predicted by circuit simulation software when applied to a vacuum capacitor-resistor circuit. A microscopic understanding of the sum of exponentials model involves charge on the surface of the circuit components.
@@ Acknowledgements.
Two oscilloscopes (Picoscope models $4262$ and $4226$) with either $16$ bit resolution at $10$ MS/s or $12$ bit resolution at $50$ MS/s along with a digital voltmeter (Keysight 34465A with input impedance greater than $1~{}\textrm{G}\Omega$) were used to collect data. The oscilloscope probes were set on a x10101010 scale with an input resistance $10~{}\textrm{M}\Omega$ and a capacitance of $15$ pF. A Comet vacuum capacitor model CFMN-500AAC/12-DE-G was used in figs. 1, 2 (inductance $<6~{}\textrm{nH}$).
A mechanical switch coupled and decoupled the battery from the circuit. No difference in the measurements was observed between the use of toggle or micro switches. The circuit layout was similar to that shown in fig. 1. The wire diameter was $0.5$ mm. All the circuits used $1/4$ W $1\%$ metal film resistors (inductance $<2~{}\textrm{nH}$).
The disk ceramic capacitor (diameter $3\times 10^{-3}~{}\textrm{m}^{2}$ and thickness $2\times 10^{-3}~{}\textrm{m}$) allowed the area enclosed by the circuit to vary from $3\times 10^{-4}$ to $3\times 10^{-2}$ $\textrm{m}^{2}$. However, the data shown in fig. 1 (b) did not then change noticeably, demonstrating that inductance is not an important parameter in these results. The similarity in response of a circuit where $dI/dt$ varies by five orders of magnitude is additional evidence against inductance being a important variable.
@@
Figure 5: Percent difference of the current between the data and exponential decay for the circuit used in fig. 2b. The upper and lower frames are magnitude and phase graphs.
Figure 6: Frequency domain results for the circuit used in figs. 3b presented in the same manner as that in fig. 5.
@@
To better illustrate the real and ideal behavior in the frequency domain the difference between the currents in the real and ideal circuits is first determined as a function of frequency. This difference when divided by the ideal current at each frequency yields a percent difference. The magnitude and phase of these percent differences are presented in fig. 5 for the circuit used in fig.2 with $R_{1}=7850~{}\Omega$ and $C=505$ pF. Results from this same procedure for the $R_{1}=59.5~{}\mathrm{K}\Omega$ and $C=1~{}\mu$F circuit used to collect the data shown in fig. 3b are given in fig. 6 .

### 01 Mar 2024 v5

# Energy relaxation in a vacuum capacitor-resistor circuit: measurement of multiple decays with divergent time constants

@@ Abstract
The decay of the electrical energy in a resistor-vacuum capacitor circuit is shown to involve multiple relaxation processes, with dramatically different time constants. This is measured using a vacuum capacitor to eliminate the effect of a dielectric between the plates (polypropylene capacitors are shown to exhibit similar behavior). A simple phenomenological model accounts for this behavior in spite of the difficulty in applying Maxwell’s equations to such a circuit. These results will lead to a revision of our understanding of the physics of circuits, having particular impact on applications that use capacitors as sensors in collecting precision data (such as found in quantum measurements and dielectric spectroscopy).
@@ I Introduction
To accurately model decay in a vacuum capacitor one must turn to Maxwell’s equations. However, only a handful of solutions that involve steady state behavior have been addressed analytically. Muller (2012) These require a surface charge on the circuit wires to maintain the current as described theoretically Sommerfeld (1952); Heald (1984); Chabay and Sherwood (2019) and utilized experimentally (where the surface charge functions as a Hall probe). Schade et al. (2019) Applying Maxwell’s equations numerically to the relaxation of an RC circuit yields an initial non-exponential decay due to transit time effects. Preyer (2002) However, this calculation does not predict the other decays that are documented below.
@@
There are few first principles analytical models of transient behavior in circuits, even obtaining numerical solutions is challenging. Klee (2020); Muller (2012); Heald (1984) The objective here is to provide a phenomenological model for the observed novel circuit behavior and relate that to a microscopic description of this energy relaxation process.
@@ II Results
Figure 1: Data for the circuit in the inset of frame (a). The upper lines in frame (a) illustrate currents through $R_{1}$ while the lowest line corresponds to the currents through $R_{2}$ for different vacuum capacitor values while $R_{1}=R_{2}=337.8~{}{\textrm{k}}\Omega$. Frames (c) and (d) show currents through $R_{1}$ (solid lines) and $R_{2}$ (dashed lines) using polypropylene capacitors. In (c) the three line pairs are for fixed $R_{1}=R_{2}=121~{}{\textrm{k}}\Omega$ while the capacitance varies from top to bottom. The lower two pairs of lines are offset along the vertical axis for clarity. The dots in the lowest trace of frame (c) are a fit using a sum of two exponential model to the $2012$ nF $R_{1}$ data. The parameters of this model for the vacuum capacitor are shown in frame (b) where the line connecting the data is only used for clarity. In (d) the capacitance is fixed at $2012$ nF while the resistance increases for the line pairs from top to bottom. The lower three line pairs have been offset from the upper pair along the vertical axis to prevent overlap. A $98$ V power supply was used to energize the circuit.
Relaxation plots of current vs. time in time constant units are shown in figs.1 (a) for vacuum capacitors and in figs.1 (c) and (d) for polypropylene capacitors. The circuit used is in the inset of fig. 1 (a). $M_{1}$ and $M_{2}$ represent the meters that measure the voltages and by Ohm’s law these currents.
Fig.1 (a) uses $R_{1}=R_{2}=337.8~{}{\textrm{k}}\Omega$ while the capacitance varies. The data for the current through $R_{1}$ in this frame are given by the upper two traces that almost overlap and have C$=0.5177$ nF and C$=1.0354$ nF respectively (the circuit capacitance is constructed from a single $0.5177$ nF capacitor and a parallel combination of two single $0.5177$ nF capacitors). Below these are two $R_{1}$ current data traces that overlap and have C$=2.789$ nF and C$=5.578$ nF respectively (constructed from a single $2.789$ nF capacitor and a parallel combination of two single $2.789$ nF capacitors). All the $R_{2}$ current data for fig.1 (a) overlap forming the lowest trace in this fig. (demonstrating single RC exponential decay down to the noise baseline).
The three pairs of solid and dashed lines in fig. 1 (c) illustrate the decay current through $R_{1}=121.5~{}\textrm{k}\Omega$ (solid line) and the decay current through $R_{2}=121.8~{}\textrm{k}\Omega$ (dashed line) for the circuit shown in the inset of fig. 1 (a). The resistance is constant while the polypropylene capacitor values of $6.947$ nF, $475.0$ nF, and $2012$ nF increase from top to bottom. The two lower line pairs are offset along the vertical axis to prevent overlap. Fig. 1 (d) shows the data for a fixed capacitance of $2012$ nF while $R_{1}=R_{2}=47.0~{}{\textrm{k}}\Omega,121~{}{\textrm{k}}\Omega,180~{}\textrm{ k}\Omega,~{}\textrm{and}~{}432~{}\textrm{k}\Omega$ increases from top to bottom. All line pairs except the pair for $R_{1}=R_{2}=47.0~{}\textrm{k}\Omega$ are offset along the vertical axis to prevent overlap.
The lowest line in fig. 1 (c) is a reproduction of the $R_{1}=R_{2}=121~{}\textrm{k}\Omega$ data using the $2012$ nF polypropylene capacitor. The filled circles that overlap with this line (characteristic of a typical fit of the data) are generated from a sum of two exponentials model, the first term of which corresponds to the expected exponential decay for an RC circuit. The parameters of the second term are shown in fig. 2 and described below. These parameters for the vacuum capacitor are shown in fig. 1 (b). Careful examination of this fit, particularly for longer times (not shown), demonstrate that the two exponential model is inadequate. However, for the majority of the data presented here this model is supported by the data within error.
@@
The unequal currents into and out of the capacitors shown in fig. 1 result in a net charge on the capacitor. However, the initial charge on each capacitor plate is not measured. In analogy with the net charge on a hairpin wire this net charge plays a role in the decay behavior.
Figure 2: The decay data from figs. 1 (c) and 1 (d) are fit with a sum of two exponentials model. The decay constant of the smaller exponential term is shown in frames (a) and (c) while the coefficient of this exponential is given in frames (b) and (d). Frames (c) and (d) use the fig. 1(c) data while frames (a) and (b) use the fig. 1(d) data. To prevent overlap of the data in these figures the data associated with $R_{2}$ are offset to the right of the $R_{1}$ data along the horizontal axes in all frames. The line connecting the data points is only used for clarity.
@@ III Phenomenological Model
A model for the fig. 1 data is given by $I=I_{10}\exp[-\alpha t]+I_{20}\exp[-\beta t]$. This is referred to as the sum of exponentials model. For all of the fits to the data presented here $\alpha=1/\textrm{RC}$, where the values of R$=R_{1}+R_{2}$ and C are determined from measurements of the individual circuit components.
The parameters of this model associated with the data of fig. 1 (c) and (d) are shown in fig. 2. Fig. 1(c), which varies C for fixed $R_{1}=121~{}\textrm{k}\Omega$, has the expected model parameters $I_{10}=0.405$ mA and $\alpha=1650,1180,17.30,4.133$ (from smaller to larger values of C) while the values of $I_{20}$ are shown in fig. 2(d) and the values of $\beta$ are shown in fig. 2(c). Fig. 1(d), which varies R for fixed C$=2012$ nF has the expected model parameters $I_{10}=1.042~{}\textrm{mA},0.4049~{}\textrm{mA},0.2722~{}\textrm{mA},0.1166~{}\textrm{mA}$ and $\alpha=10.63,4.132,2.777,1.190$ (from smaller to larger values of R) while the values of $I_{20}$ are shown in fig. 2(b) and the values of $\beta$ are shown in fig. 2(a).
The model parameters for current flowing through $R_{1}$ in the circuit of fig. 1 (a) using a fixed vacuum capacitor with varying resistance is similar to that shown in fig. 2 (a) and (b). The similarity in the decay behavior of the vacuum and polypropylene capacitors is manifest in these model parameters as documented in figs. 1 (c) and 2. This suggests that the decay mechanisms observed are intrinsic to capacitors.
Plots of current vs time in time constant units allow data to be presented over a range of RC values. However, the parameters of the sum of two exponentials model cannot be inferred from such graphs and must be fit to the current vs time data.
To measure the decay as the charge distributes over an isolated conductor the single capacitor in fig. 1 is replaced with two in series, as shown in the inset of fig 3 (a) with switches that operate synchronously (within a millisecond while the time constant is $60$ milliseconds). When the switches are closed the isolated right plates of polypropylene capacitors $C_{1}=2.012~{}\mu F$ and $C_{2}=2.014~{}\mu F$ have charge and no charge, respectively. When the switches open the right plates then have a initial net charge which is distributed between them during the decay. Without a switch across $C_{2}$ the net charge on the right plates of $C_{1}$ and $C_{2}$ is initially zero (they are grounded before data collection commences). The response of the circuit for these two cases is shown in figs. 3 (b) and (c). The latter demonstrates enhancement in the non-exponential decay due to this net charge on the right plates of $C_{1}$ and $C_{2}$.
@@
Figure 3: Frames (a), (c), and (d) show meter $M_{2}$ voltage data for the circuit in the inset of frame (a). Voltage data for meter $M_{1}$ is shown in frame (b). The insets in frames (b) and (c) show $C_{dyn}$ for the respective models of the $M_{2}$ data (in microfarads for (b)) while (d) shows the response for different initial voltages. The capacitors are each $2.01~{}\mu F$. For frame (a) through (c) $R=59.5~{}\mathrm{k}\Omega$ with $V=9.8$. For frame (d) $R=578~{}\mathrm{k}\Omega$ with $V=20,0.74$, and $0.088$ with a microvolt noise baseline and error bars displayed only for a few of the data points.
@@ IV Dynamic Capacitance Model
Figure 4: The decay measured for $V_{1}=10$ with different initial voltages on the right side of the polypropylene capacitor for the circuit in frame (a) with $R=274~{}\mathrm{k}\Omega$ and $C=2.01~{}\mu$F. The switches are closed to energize the capacitor. They then open simultaneously for the discharge. Frames (b) and (c) illustrate decays for $V_{2}=0,6,9,9.8$ and $V_{2}=20,14,11.5,10.15$ from top to bottom, respectively. The time constants in seconds for the $V_{20}\exp[-\beta t]$ (plotted with error bars) and $V_{10}\exp[-\alpha t]$ (plotted with filled circles) decays are shown in frame (d).
@@ V Surface Charge Model
Another modeling approach is to separate the amount of charge on a capacitor plate into two parts: one associated with the energy stored in between the capacitor plates, $Q^{in}$, whose electric field is confined in between the plates, and $Q^{out}$, the charge on the circuit wire and capacitor plate that facilitates the flow of current in the circuit with its electric field not confined. The superposition principle yields a response that is a sum of responses for each type of charge in this separate charge model. Let the $V_{10}\exp[-\alpha t]$ term in the above sum of exponentials model be associated with $Q^{in}$. Let $Q^{out}_{0}$ be the charge required to direct the current from the power supply into $R$ before the switch is opened. This is expressed as $Q^{out}_{0}=V^{out}_{0}\tau_{out}/R$, where $\tau_{out}$ is a constant and $V^{out}_{0}$ is the voltage at the plate generated by $Q^{out}_{0}$. For smaller $R$ larger $Q^{out}_{0}$ is required to direct this larger current away from the capacitor plate and into the resistor (similar to the effect that charge on the hairpin wire has on the current flowing through the wire).
@@ VI Discussion
The geometry of a circuit increases the difficulty in calculating $Q^{out}$ from Maxwell’s equations. Muller (2012) Even constructing a numerical model that predicts only a single exponential decay for an RC circuit from Maxwell’s equations is nontrivial. Preyer (2002) Yet the data illustrate similar but simple (a sum of exponentials) behavior for geometrically dissimilar capacitors. An additional issue in solving Maxwell’s equations is a potential coupling between these decays. However, currents due to $Q^{out}$ are measured even after $Q^{in}$ has exponentially decayed over tens of time constants. During such times $Q^{in}$ is so small that coupling between these decays should not be significant.
@@ VII Conclusions
The decay of the electrical energy in resistor-vacuum capacitor and resistor-polypropylene capacitor circuits involves multiple relaxation processes with divergent time constants and net charge on the capacitor. This novel behavior differs fundamentally from the single exponential decay predicted by circuit simulation software. A microscopic understanding of the sum of exponentials model, that the data support, involves charge on the surface of the circuit components.
@@ Acknowledgements.
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@
A Keysight 34465A digital voltmeter was used to collect data with input impedances of $10^{7}~{}\Omega$ for the data in fig. 1 and greater than $10^{9}~{}\Omega$ for the data in fig. 3. At least fifteen samples were taken for the vacuum capacitor data and collected at a $50$ kHz rate. The data was process by first taking a mean value of these samples. Next a moving average of four to twenty points was performed on the logarithm of the mean of these samples. One power line cycle integration during the data acquisition was applied to the longer time constant data ($2\mu$F polypropylene capacitor circuits).
A mechanical switch coupled and decoupled the power supply from the circuit (with an open switch capacitance of less than $0.1$ pF). The circuit layout was similar to that shown in fig. 1. The wire diameter was $0.5$ mm. All the circuits used $1/4$ W $1\%$ metal film resistors (inductance $<2~{}\textrm{nH}$). Comet vacuum capacitors models CFMN-2800BAC/8-DE-G and CFMN-500AAC/12-DE-G were used (inductance $<6~{}\textrm{nH}$). The circuit was enclosed in a shielded box.
The small $7$ nF polypropylene capacitor allowed the area enclosed by the circuit to vary from $3\times 10^{-4}$ to $3\times 10^{-2}$ $\textrm{m}^{2}$. However, the decay data shown in fig. 1 (b) did not then change noticeably, demonstrating that inductance is not an important parameter in these results. The similarity in response of a circuit where $dI/dt$ varies by five orders of magnitude is additional evidence against inductance being a important variable.
@@
Figure 5: Frequency domain results are shown for the circuit in figs. 3b. Percent differences of the current between the data and exponential decay are given. The upper and lower frames are magnitude and phase graphs.
@@
To better illustrate the real and ideal behavior in the frequency domain the difference between the currents in the real and ideal circuits is first determined as a function of frequency. This difference when divided by the ideal current at each frequency yields a percent difference. The magnitude and phase of these percent differences are presented in fig. 5 for the circuit used in fig.3.